The graph of y=ax^2 + bx + C opens up if a > 0 and down if a < 0. Explain why or how that staement is true. What happens as x gets more a more positive or more and more negative?
Answer by stanbon(75887)

The graph of y=ax^2 + bx + C opens up if a > 0 and down if a < 0. Explain why or how that staement is true.
--------------------------------------
What happens as x gets more a more positive or more and more negative?
------------------------------------
1st: You have to understand that the ax^2 overwhelms the effect of bx+c
as x gets more and more positive or gets more and more negative.
----
2nd: So if a>0, y = ax^2 gets higher and higher if x gets more and more
positive OR if x gets more and more negative.
----
And if a<0, y = ax^2 gets more and more negative if x gets more and more
positive OR if x gets more and more negative.
------------------------------------------------
Cheers,
Stan H.